Please help me solve this. The length of a rectangular lot is 50 feet more than the width. If the perimeter is 500 feet, then what are the length and width?
Answer by ankor@dixie-net.com(22740)

The length of a rectangular lot is 50 feet more than the width.
If the perimeter is 500 feet, then what are the length and width?
:
2L + 2W = 500
Simplify, divide by 2:
L + W = 250
:
"The length of a rectangular lot is 50 feet more than the width."
L = W+50
:
Replace L with (W+50) in the simplified equation; L + W = 250
(W+50) + W = 250
2W = 250 - 50
2W = 200
W = 100 is the width
and
100+50 = 150 is the length
:
:
Check solution by finding the perimeter
2(150) + 2(100) = 500